OpenStudy (iwanttogotostanford):

What is the simplified form of the quantity of x plus 9, all over the quantity of 8 − the quantity of x plus 3, all over the quantity of x plus 2? the quantity of x squared plus 19x plus 42, all over 8 times the quantity of x plus 2 the quantity of x plus 6, all over the quantity of 6 minus x 6 over the quantity of 6 minus x the quantity of x squared plus 3x minus 6, all over 8 times the quantity of x plus 2

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo

OpenStudy (anonymous):

I presume that phrasing to be read as: \[\frac{ x + 9 }{ 8 } - \frac{ x + 3 }{ x + 2 }\] If this is the case, then: \[\frac{ x + 9\left( x + 2 \right) - 8\left( x + 3 \right) }{ 8\left( x + 2 \right) }\] When simplified it becomes: \[\frac{ x ^{2} + 3x - 6 }{ 8(x + 2) }\] How would this be worded? @iwanttogotostanford

OpenStudy (mathmale):

How was this problem presented in your homework assignment? If in symbols, I'd strongly suggest that you learn to present a function in either or both the Equation Editor or the Draw utility (below). It's additional work for helpers if y ou present the problem statement in words, as well as an invitation to error.